Opened 10 years ago

Closed 8 years ago

#497 closed flight processing (fixed)

IG13/21, flight day 222b/2013, Helheim

Reported by: knpa Owned by: benj
Priority: alpha 4 high Milestone: 2013 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by tipo)

Data location: /users/rsg/arsf/arsf_data/2013/flight_data/greenland_iceland/IG13_21-2013_222b_Helheim/

Data arrived from ARSF via network transfer on 23/08/2013

IG13/21 Scientific objective: Testing a theory of glacial calving.

IG13/21 Priority: 9

PI: Tavi Murray

Sensors:

Camera (Delivered 07/10/2013)
Eagle (Delivered 07/10/2013)
Hawk (Delivered 07/10/2013)
LiDAR (Delivered 07/10/2013)

Change History (17)

comment:1 Changed 10 years ago by stgo

Processing nav using temporary basestation, will reprocess hyperspectral and RCD once new basestation data comes in. Lidar will wait until new basestation data arrives.

comment:2 Changed 10 years ago by stgo

Hyperspectral SCTs:

Flightline Eagle SCT Hawk SCT
1 -0.01 -0.01
2 -0.01 -0.01
3 -0.01 -0.01
4 -0.01 -0.01
5 -0.01 -0.01
6 -0.01 -0.01
7 -0.00 -0.01
8 -0.01 -0.01
9 -0.01 -0.01
10 -0.01 -0.01

comment:3 Changed 10 years ago by stgo

Processing using new basestation data. PPP came back with quality 2 throughout.
PPP Result:

Lat 66 24 02.38098
Long -38 12 50.55486
Height 421.240

comment:4 Changed 10 years ago by stgo

Processed nav, separation and accuracy results are both good.

comment:5 Changed 10 years ago by stgo

Reprocessing hyperspectral, beginning lidar processing.

comment:6 Changed 10 years ago by stgo

Line 9 was aborted according to logsheet, presumably why it has 37% returns and is so short, won't be included in delivery.

comment:7 Changed 10 years ago by tipo

LiDAR delivery checked, no problems found.

Version 0, edited 10 years ago by tipo (next)

comment:8 Changed 10 years ago by stgo

Hyperspectral ready for delivery check. Updated scts:

Flightline Eagle SCT Hawk SCT
1 -0.01 -0.01
2 -0.01 -0.01
3 -0.04 0.00
4 -0.01 -0.13
5 -0.01 -0.01
6 -0.01 -0.01
7 0.00 -0.01
8 -0.01 -0.01
9 -0.01 -0.01
10 -0.01 -0.01

comment:9 Changed 10 years ago by tipo

LiDAR delivery checked, no problems found.

comment:10 Changed 10 years ago by tipo

Hyperspectral delivery checked, no problems found.

comment:11 Changed 10 years ago by stgo

Camera ready for delivery check.

Last edited 10 years ago by stgo (previous) (diff)

comment:12 Changed 10 years ago by stgo

Changed camera readme according to tipos comments.

comment:13 Changed 10 years ago by tipo

  • Component changed from Processing: general to Archiving
  • Description modified (diff)
  • Owner set to benj

comment:14 Changed 10 years ago by tipo

Dispatched deliveries 07/10/2013

comment:15 Changed 8 years ago by dap

Starting archiving

comment:16 Changed 8 years ago by dap

Archiving

rsync to NEODC complete

comment:17 Changed 8 years ago by dac

  • Resolution set to fixed
  • Status changed from new to closed

Archiving

Data now available from NEODC.

Note: See TracTickets for help on using tickets.